# Pierre Dantas Dept. of Computer Science

### v1

@@ Abstract
The idea of using Formal Verification tools with large language models (LLMs) has enabled scaling software verification beyond manual workflows. However, current methods remain unreliable. Without a solid theoretical footing, the refinement process can wander; sometimes it settles, sometimes it loops back, and sometimes it breaks away from any stable trajectory. This work bridges this critical gap by developing an LLM-Verifier Convergence Theorem, providing the first formal framework with provable guarantees for termination and convergence. We model the interaction between the LLM and the verifier as a discrete-time Markov Chain, with state transitions determined by a key parameter: the error-reduction probability ($\delta$). The procedure reaching the Verified state almost surely demonstrates that the program terminates for any $\delta>0$, with an expected iteration count bounded by $\mathbb{E}[n]\leq 4/\delta$. We then stress-tested this prediction in an extensive empirical campaign comprising more than 90,000 trials. The empirical results match the theory with striking consistency. Every single run reached verification, and the convergence factor clustered tightly around $C_{f}\approx$ 1.0. Consequently, the bound mirrors the system’s actual behavior. The evidence is sufficiently robust to support dividing the workflow into three distinct operating zones: marginal, practical, and high-performance. Consequently, we establish the design thresholds with absolute confidence. Together, the theoretical guarantee and the experimental evidence provide a clearer architectural foundation for LLM-assisted verification. Heuristic tuning no longer has to be carried out by the system. Engineers gain a framework that supports predictable resource planning and performance budgeting, precisely what is needed before deploying these pipelines into safety-critical software environments.
Keywords Formal Verification $\cdot$ Large Language Models $\cdot$ SMT Solvers $\cdot$ Bounded Model Checking $\cdot$ Automated Program Repair $\cdot$ Specification Synthesis $\cdot$ ESBMC
@@ 1 Introduction
The Fundamental Gap addressed in this work: Unpredictable Artificial Intelligence (AI)-Verifier Interactions. We cannot predict how AI and Verifiers will interact. Mixing statistical AI components with reliable, deterministic verification tools creates a core conflict that current methods have not solved mathematically. Even though specific tests look promising [Pirzada2024LLM, Lin2024FVEL], we do not have a reliable convergence theory for those repeated refinement cycles, which is now a trend in the field. This missing piece creates three problems that stop us from using these systems in real-world applications:
@@
    Unpredictable Termination: Without mathematical guarantees that it will converge, the refinement process can get stuck in a loop, endlessly oscillating between states, or wander off track. This unpredictability makes it unreliable when applied to scenarios such as flight control or medical systems, where a predictable, guaranteed end to the process is mandatory.
@@
    Unbounded Resource Consumption: The absence of a defined step limit precludes the configuration of reliable timeouts or the accurate allocation of computing resources. In a real-world system, this is a significant risk – the process could run out of control, consuming all available memory or processing time until the entire application crashes.
@@
    Lack of Performance Predictability: It is currently a guessing game. Engineers have no systematic way to connect a LLMs’ capabilities to its automated verification performance, forcing them to make architectural decisions through trial and error rather than on solid mathematical principles.
These limitations can lead to serious problems in the real world. In particular, they can make systems less safe, waste computer resources, create challenges for getting approved, and make it hard for engineers to make important design decisions without enough information.
Our Proposed Approach: A Convergence Theorem for LLM-Verification. We tackle the problem of guarantees using a new Convergence Theorem. We found that even though the LLM has some randomness, the way we improve it follows a pattern such as a Markov Chain. By seeing it this way, we can show that it will definitely come to an end, rather than relying on uncertain guesses [Zekri2024, Jianing2024].
Simply put, $\delta$ quantifies an LLM’s single-attempt success rate for a verification task. This parameter gives important information to help understand how the model works. It allows users to feel confident using the system in serious engineering projects. The value of $\delta$ is not only a theory; real measurements from actual tasks show that it is useful [Pirzada2024LLM]. For example, it ranges from $0.35$ for making things that stay the same to $0.65$ for creating proofs [First2023Baldur].
@@
  • If the program terminates and all reachable states are checked, the process is guaranteed to eventually reach the Verified state, regardless of its initial state.
  • The process is efficient. We can guarantee that the average number of steps needed will never exceed ${4}/{\delta}$, providing a worst-case ceiling for resource planning.
  • The likelihood of the process needing more than $k$ steps drops off exponentially. This sharp decline enables reliable timeouts with high confidence.
Together, these results directly address the core challenges we identified here: our framework guarantees the process will finish, imposes a hard limit on computational costs, and provides engineers with a mathematical basis, via the $\delta$ parameter, for predicting system performance.
To confirm our framework, we conducted an extensive study involving more than 90,000 simulations spanning the range 0.1 to 0.9. Our experimental results strongly support our theoretical predictions and show three distinct behavior patterns. Each of these patterns has significant implications for how real-world systems should be designed (see Table 1).
@@
Marginal $\delta<0.3$ Suitable for applications where performance requirements are relaxed.
Practical $0.3\leq\delta\leq 0.6$ Optimal for real-world verification tasks with balanced performance.
High-Performance $\delta>0.6$ Ideal for safety-critical systems requiring fast convergence.
@@ 3.1.3 Markovian Properties of LLM-Verifier Interactions
In our framework, the system reaches the Verified state as its final state. The Unverified and CandidateGenerated states are temporary states that the system can pass through before reaching the final state. Recent studies show that LLMs naturally follow these state changes [Zekri2024, Jianing2024], which matches well with Formal Methods that describe “verification” as a process moving through different states [Clarke1997, Clarke2018Handbook].
@@ 4 Proposed Work - LLM-Verifier Convergence Theorem
The main goal of combining LLMs with Formal Verification tools like ESBMC is simple: Can we prove with math that repeated improvements will finally lead to a correct answer? To answer this, we build upon the Markov Chain theory (detailed in Section 3.1.3) to develop a new Theorem tailored for LLM-verifier systems.
@@ 4.1 Theoretical Foundation and Markov Chain Modeling
Eq. (5) (p. 5) displays the canonical form of the transition matrix $P$. Here, the fundamental matrix $N=(I_{t}-Q)^{-1}$ gives the expected number of visits to transient states before absorption. We then use the expected time to absorption from Eq. (6) (p. 6) to build our convergence analysis.
Our model treats the LLM-verifier interaction as an absorbing Markov Chain, as outlined in Section 3.1.3. This approach is well-supported theoretically and practically. Since transformer-based LLMs generate text token-by-token within limited context windows, they inherently display the properties of a Markov process. At the same time, empirical studies of systems like ESBMC-AI [Tihanyi2025New] and Baldur [First2023Baldur] confirm that transition probabilities mainly depend on the current verification state.
This model uses a five-state Markov Chain with Verified as the absorbing state. In this model, we treat the refinement steps - Unverified and CandidateGeneratedi - as transient states. The logic here mimics an actual iterative loop. When the system is in a CandidateGenerated state, it attempts to move to Verified. The transition succeeds with probability $\delta$. If it fails (with probability $1-\delta$), the system stays in the loop to try again.
The Error-Reduction Probability ($\delta$): Central to our model is the error-reduction probability $\delta$, which quantifies the LLM’s capability to generate correct verification artifacts in a single attempt. Table 5 shows typical $\delta$ values (0.360 - 0.880) for LLM verification tasks, validating its role in governing system convergence and providing the empirical foundation for our theory.
@@ 4.2 The Main Result: LLM-Verifier Convergence Theorem
We model the LLM-verifier process using a discrete-time Markov Chain, denoted as $X=\{X_{n}\}_{n\geq 0}$. The state space $S$ consists of two parts ($S=T\cup A$). First, $T=\{s_{1},s_{2},s_{3},s_{4}\}$ represents the set of transient refinement attempts. Second, $A=\{s_{5}\}$ acts as the single absorbing state. Assuming a fixed success probability $\delta\in(0,1]$, we construct the transition matrix $P=(P_{i,j})$. The individual entries $P_{i,j}=\mathbb{P}(X_{n+1}=s_{j}\mid X_{n}=s_{i})$ are arranged as follows:
@@
    Transient States: For the transient states $i\in\{1,2,3,4\}$, the transition probabilities are:
    $P_{i,j}=\begin{cases}\delta&\text{if }j=i+1\quad\text{(progress to next state),}\\ 1-\delta&\text{if }j=i\quad\text{(retry current step),}\\ 0&\text{otherwise.}\end{cases}$
@@
    Absorbing State: State $s_{5}$ is absorbing, meaning the process stops here. Hence, $P_{5,5}=1$.
@@
Part 2: Expected Iteration Bound: We can analyze the total iteration count $\tau$ by considering the time spent in each state $s_{j}$. Let $M_{j}$ represent the number of steps the process remains in state $s_{j}$. Because the probability of exiting $s_{j}$ on any given step is $\delta$, the sojourn time $M_{j}$ follows a geometric distribution, $M_{j}\sim\text{Geom}(\delta)$, with an expected value of $\mathbb{E}[M_{j}]=1/\delta$. Since the total number of iterations is simply the sum of the time spent in all states, $\tau=\sum_{j=1}^{4}M_{j}$, the linearity of expectation immediately gives us:
$\mathbb{E}[\tau\mid X_{0}=s_{1}]=\sum_{j=1}^{4}\mathbb{E}[M_{j}]=\frac{4}{\delta}$
Part 3: Tail Bound: The transient submatrix $Q$ is diagonal. Specifically, every diagonal entry is $1-\delta$. As a result, the spectral radius is $\lambda_{Q}=1-\delta$. Using standard Markov Chain theory, we find:
@@ 4.3 A Practical Framework for Convergence
The system’s behavior is governed by the error-reduction probability $\delta$, a metric that quantifies the LLM’s capability to generate correct verification artifacts. From any CandidateGenerated state, the process has a $\delta$ chance of successfully advancing to Verified; otherwise, with a probability of $1-\delta$, it persists in the refinement loop.
Figure 4 illustrates our integrated LLM-Verifier Convergence Framework, which combines theoretical modeling, convergence analysis, and empirical validation. Our methodology focuses on building hybrid verification systems. We model the refinement steps as a Markov Chain. This approach yields theoretical guarantees, which we further validate through practical simulation. A feedback loop links real-world testing to the theoretical model, allowing us to adjust parameters - especially the error-reduction probability $\delta$ - and improve the model’s assumptions step by step.
Figure 4: LLM-Verifier Convergence Framework integrating theoretical modeling, convergence analysis, and empirical validation
Our theory represents an important advancement in bringing together LLMs and Formal Methods. We no longer have to depend on uncertain outcomes. Instead, we offer a straightforward and dependable framework that is backed by mathematical evidence. Our theory offers a significant step forward for combining LLMs and Formal Methods. We no longer need to rely on uncertainty. In its place, we provide a clear, reliable framework supported by mathematical proofs.
  • Resource Allocation: Provides helpful guidelines, like $\mathbb{E}[n]={4}/{\delta}$, to assist with setting time limits and planning how much computing power to use.
  • Capability Assessment: Introduces a clear metric ($\delta$) to measure LLM performance on verification tasks.
  • Deployment Guidelines: Establishes practical thresholds for use ($\delta>0.3$ for standard systems and $\delta>0.6$ for high-performance applications.applications).
Mathematically, we confirm that for any $\delta>0$, the system will reach the confirmed state. We can then use the formula $\mathbb{E}[n]={4}/{\delta}$ to plan our system limits, which converts a theoretical measure into a usable engineering metric. Table 6 shows how to use this logic to distribute resources effectively in various situations.
Table 6: Expected iteration bounds for different LLM verification capabilities
LLM Capability Level Success Rate $\delta$ Expected Iterations $\mathbb{E}[n]$
@@
Furthermore, the exponential tail bound $\mathbb{P}(n>k)\leq\alpha(1-\delta)^{k}$ ensures that the probability of excessively long refinement loops decays exponentially, enabling reliable timeout configurations for safety-critical systems. This mathematical framework bridges theoretical guarantees with practical deployment considerations, providing both foundational rigor and actionable engineering insights.

### v2

@@ Abstract
The integration of Formal Verification tools with Large Language Models (LLMs) offers a path to scale software verification beyond manual workflows. However, current methods remain unreliable: without a solid theoretical footing, the refinement process acts as a black box that may oscillate, loop, or diverge. This work bridges this critical gap by developing an LLM-Verifier Convergence Theorem, providing the first formal framework with provable guarantees for termination in multi-stage verification pipelines. We model the interaction not as a generic loop, but as a sequential absorbing Markov Chain comprising four essential engineering stages: CodeGen, Compilation, InvariantSynth, and SMTSolving. We prove that for any non-zero stage success probability ($\delta>0$), the system reaches the Verified state almost surely. Furthermore, because of the sequential nature of the pipeline, we derive a precise latency bound of $\mathbb{E}[n]\leq 4/\delta$. We stress-tested this prediction in an extensive empirical campaign comprising over 90,000 trials. The results match the theory with striking consistency: every run reached verification, and the empirical convergence factor clustered tightly around $C_{f}\approx 1.0$, confirming that the $4/\delta$ bound accurately mirrors system behavior rather than serving as a loose buffer. Based on this data, we identify three distinct operating zones – marginal, practical, and high-performance – and propose a dynamic calibration strategy to handle parameter drift in real-world environments. Together, these contributions replace heuristic guesswork with a rigorous architectural foundation, enabling predictable resource planning and performance budgeting for safety-critical software.
keywords:
Formal Verification, Large Language Models, SMT Solvers, Bounded Model Checking, Automated Program Repair, Specification Synthesis, ESBMC
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
@@ 1 Introduction
The Fundamental Gap addressed in this work: Unpredictable Artificial Intelligence (AI)-Verifier Interactions. We cannot predict how AI and Verifiers will interact in multi-stage workflows. Mixing statistical AI components with reliable, deterministic verification tools creates a core conflict that current methods have not solved mathematically. Even though specific tests look promising [Pirzada2024LLM, Lin2024FVEL], we lack a reliable convergence theory for the repeated refinement cycles typical of modern verification pipelines. Most existing approaches treat LLM refinement as a generic “black box” loop, ignoring the sequential engineering dependencies required to build valid proofs. This missing piece creates three problems that prevent the deployment of these systems in real-world applications:
@@
    Unpredictable Termination: Without mathematical guarantees that the sequential process will converge, the refinement can get stuck in intermediate stages (e.g., oscillating between syntax errors and logical failures) or wander off track. This unpredictability makes it unreliable for safety-critical scenarios like flight control, where a guaranteed end to the process is mandatory.
@@
    Unbounded Resource Consumption: The absence of a defined step limit precludes the configuration of reliable timeouts. In a multi-stage pipeline, costs accumulate; without a theoretical ceiling, the process could run out of control, consuming all available memory or processing time until the application crashes.
@@
    Lack of Performance Predictability and Stability: It is currently a guessing game. Engineers have no systematic way to connect an LLM’s capability to its total verification latency. Furthermore, the assumption that LLM performance is constant (stationary) is often violated in practice, leaving engineers without strategies to detect or mitigate performance drift during operation.
These limitations can lead to serious problems in the real world, making systems unsafe and rendering resource planning impossible.
Our Proposed Approach: A Convergence Theorem for Sequential LLM-Verification. We tackle the problem of guarantees using a new Convergence Theorem based on a sequential absorbing Markov Chain. We model the verification process not as a generic loop but as a structured pipeline comprising four essential engineering stages: CodeGen, Compilation, InvariantSynth, and SMTSolving.
Figure 1 illustrates our integrated LLM-Verifier Convergence Framework. The diagram depicts the sequential flow from the initial Unverified state through four milestones – CodeGen ($s_{1}$), Compilation ($s_{2}$), InvariantSynth ($s_{3}$), and SMTSolving ($s_{4}$) – to the final Verified absorbing state ($s_{5}$). By modeling these dependencies as a sequential absorbing Markov Chain, where $\delta$ governs the transition to the next stage ($s_{i}\to s_{i+1}$), we derive rigorous theoretical guarantees ($\mathbb{E}[n]\leq 4/\delta$). A feedback loop links real-world testing to the theoretical model, supporting the dynamic calibration strategy to handle parameter non-stationarity in deployment. By continuously monitoring the empirical success rate ($\hat{\delta}$) against the theoretical stability regions, this mechanism detects performance drift (e.g., when $\hat{\delta}$ drops into the marginal region) and triggers corrective actions – such as context resets or temperature adjustments – to restore the system to the optimal practical region.
Figure 1: LLM-Verifier Convergence Framework. The model visualizes the sequential progression through four mandatory refinement stages ($s_{1}\to s_{4}$), where $\delta$ represents the probability of advancing to the next stage and $1-\delta$ represents the probability of retrying the current stage. A feedback loop links real-world testing to the theoretical model, supporting the dynamic calibration strategy to handle parameter non-stationarity in deployment. By continuously monitoring the empirical success rate ($\hat{\delta}$) against the theoretical stability regions, this mechanism detects performance drift (e.g., when $\hat{\delta}$ drops into the marginal region) and triggers corrective actions – such as context resets or temperature adjustments – to restore the system to the optimal practical region
In this model, $\delta$ quantifies the LLM’s single-attempt success rate for progressing from one stage to the next. A failure (probability $1-\delta$) necessitates a retry of the current stage, while success moves the artifact closer to the final Verified state. This structured view allows us to move beyond uncertain guesses [Zekri2024, Jianing2024] and prove that the system will definitely terminate.
@@
  • Guaranteed Convergence: If the program terminates and all reachable states are checked, the process is guaranteed to eventually traverse all four stages and reach the Verified state almost surely.
  • Bounded Latency: The process is efficient. Because the system must pass through four distinct stages ($s_{1}\to s_{4}$), we prove that the expected number of iterations is the sum of the expected time for each stage. This yields a precise bound of $\mathbb{E}[n]\leq 4/\delta$, providing a worst-case ceiling for resource planning.
  • Exponential Tail Bounds: The likelihood of the process needing more than $k$ steps drops off exponentially. This sharp decline enables reliable timeouts with high confidence.
To address the practical challenge of parameter non-stationarity (performance drift), our framework explicitly supports dynamic calibration. By continuously monitoring the empirical success rate against the theoretical $\delta$ regions, engineers can detect when an LLM drops into the marginal region and trigger context resets. Together, these results provide a comprehensive architectural foundation: our framework guarantees the process will finish, imposes a hard limit on computational costs ($4/\delta$), and provides a mechanism to maintain stability in dynamic environments.
To confirm our framework, we conducted an extensive study involving more than 90,000 simulations spanning the range 0.1 to 0.9. Our experimental results strongly support our theoretical predictions and show three distinct behavior patterns (see Table 1).
@@
Marginal $\delta<0.3$ High variance; requires strict timeouts or dynamic calibration to prevent resource exhaustion.
Practical $0.3\leq\delta\leq 0.6$ Optimal for real-world verification tasks with balanced performance and predictable latency.
High-Performance $\delta>0.6$ Ideal for safety-critical systems requiring fast convergence and minimal variance.
@@ 3.1.3 Markovian Properties of LLM-Verifier Interactions
In our framework, the system reaches the Verified state ($s_{5}$) as its final absorbing goal. The path to this goal involves traversing a sequence of distinct transient states representing engineering milestones: CodeGen ($s_{1}$), Compilation ($s_{2}$), InvariantSynth ($s_{3}$), and SMTSolving ($s_{4}$). Recent studies show that LLMs naturally follow these step-by-step reasoning patterns [Zekri2024, Jianing2024], which matches well with Formal Methods that describe “verification” as a process moving through defined refinement stages [Clarke1997, Clarke2018Handbook].
@@ 4 Proposed Work - LLM-Verifier Convergence Theorem
The main goal of combining LLMs with Formal Verification tools like ESBMC is simple: Can we prove with math that repeated improvements will finally lead to a correct answer? To answer this, we build upon the Markov Chain theory (detailed in Section 3.1.3) to develop a new Theorem tailored for sequential LLM-verifier systems.
@@ 4.1 Theoretical Foundation and Markov Chain Modeling
This approach is well-supported theoretically and practically. Since transformer-based LLMs generate text token-by-token within limited context windows, they inherently display the properties of a Markov process. At the same time, empirical studies of systems like ESBMC-AI [Tihanyi2025New] and Baldur [First2023Baldur] confirm that transition probabilities mainly depend on the current verification state.
This model uses a five-state Markov Chain with Verified as the absorbing state ($s_{5}$). Crucially, we replace generic refinement steps with four distinct transient states representing the essential milestones of a formal verification pipeline:
  1. 1.
    $(s_{1})$ CodeGen: The LLM attempts to generate the initial code or specification.
  2. 2.
    $(s_{2})$ Compilation: The artifact is validated for syntax and compilation errors.
  3. 3.
    $(s_{3})$ InvariantSynth: The system attempts to synthesize necessary loop invariants.
  4. 4.
    $(s_{4})$ SMTSolving: The final formal check where the solver attempts to prove correctness.
The logic here mimics a sequential engineering workflow. When the system is in a specific stage $s_{i}$, it attempts to advance to the next milestone $s_{i+1}$. The transition succeeds with probability $\delta$. If it fails (with probability $1-\delta$), the system remains in the current loop ($s_{i}$) to retry the specific task (e.g., regenerate code or refine invariants). The process reaches the absorbing Verified state only after successfully clearing the final stage $s_{4}$.
The Error-Reduction Probability ($\delta$): Central to our model is the error-reduction probability $\delta$, which quantifies the LLM’s capability to successfully complete a single pipeline stage in one attempt. Table 5 shows typical $\delta$ values (0.360 - 0.880) for LLM verification tasks, validating its role in governing system convergence and providing the empirical foundation for our theory.
@@ 4.2 The Main Result: LLM-Verifier Convergence Theorem
We model the LLM-verifier process using a discrete-time Markov Chain, denoted as $X=\{X_{n}\}_{n\geq 0}$. The state space $S$ consists of a sequential engineering pipeline ($S=T\cup A$). First, $T=\{s_{1},s_{2},s_{3},s_{4}\}$ represents the set of transient pipeline stages (CodeGen, Compilation, InvariantSynth, SMTSolving). Second, $A=\{s_{5}\}$ acts as the single absorbing state (Verified). Assuming a fixed success probability $\delta\in(0,1]$ for passing any single stage, we construct the transition matrix $P=(P_{i,j})$. The individual entries $P_{i,j}=\mathbb{P}(X_{n+1}=s_{j}\mid X_{n}=s_{i})$ are arranged as follows:
@@
    Transient Pipeline States: For the transient states $i\in\{1,2,3,4\}$, the transition probabilities are:
[…]
@@
    Absorbing State: State $s_{5}$ is absorbing, meaning the verification is complete. Hence, $P_{5,5}=1$.
@@
Part 2: Expected Iteration Bound: We analyze the total iteration count $\tau$ by considering the sequential time spent in each pipeline stage. Let $M_{j}$ represent the number of steps the process remains in state $s_{j}$ before advancing. Because the probability of exiting $s_{j}$ to $s_{j+1}$ on any given step is $\delta$, the sojourn time $M_{j}$ follows a geometric distribution, $M_{j}\sim\text{Geom}(\delta)$, with an expected value of $\mathbb{E}[M_{j}]=1/\delta$. Since the process is sequential, the total number of iterations is the sum of the time spent in all four stages, $\tau=\sum_{j=1}^{4}M_{j}$. The linearity of expectation immediately gives us:
$\mathbb{E}[\tau\mid X_{0}=s_{1}]=\sum_{j=1}^{4}\mathbb{E}[M_{j}]=\frac{1}{\delta}+\frac{1}{\delta}+\frac{1}{\delta}+\frac{1}{\delta}=\frac{4}{\delta}$
Part 3: Tail Bound: The transient submatrix $Q$ is upper triangular (not diagonal), as transitions only occur from $i$ to $i$ or $i+1$. Specifically, every diagonal entry is $P_{i,i}=1-\delta$, and every super-diagonal entry is $P_{i,i+1}=\delta$. For an upper triangular matrix, the eigenvalues are simply the diagonal entries. Thus, the spectral radius is $\lambda_{Q}=\max_{i}|P_{i,i}|=1-\delta$. Using standard Markov Chain theory, we find:
@@ 4.3 A Practical Framework for Convergence
The system’s behavior is governed by the error-reduction probability $\delta$, a metric that quantifies the LLM’s ability to complete a specific verification sub-task. Unlike simpler models that assume a single “generate-and-verify” step, our framework acknowledges the engineering reality of a multi-stage pipeline. From any transient pipeline stage $s_{i}$ (e.g., Compilation or InvariantSynth), the process has a $\delta$ chance of successfully advancing to the next milestone $s_{i+1}$. Conversely, with a probability of $1-\delta$, the specific stage fails, necessitating a retry ($s_{i}\to s_{i}$) within the refinement loop before the system can proceed.
Our theory represents an important advancement in bridging LLMs and Formal Methods. We no longer have to depend on uncertain, black-box outcomes. Instead, we offer a straightforward and dependable framework backed by mathematical evidence:
  • Resource Allocation: Provides precise guidelines, specifically the bound $\mathbb{E}[n]={4}/{\delta}$, to assist with setting timeout limits and planning computational budgets for the entire pipeline.
  • Capability Assessment: Introduces a clear metric ($\delta$) to measure LLM performance on specific verification tasks, allowing for modular benchmarking of different stages (e.g., measuring $\delta$ for invariant generation separately from code repair).
  • Deployment Guidelines: Establishes practical thresholds for use ($\delta>0.3$ for standard systems and $\delta>0.6$ for high-performance applications).
Mathematically, we confirm that for any $\delta>0$, the system will reach the verified state almost surely. We can then use the derived formula $\mathbb{E}[n]={4}/{\delta}$ to plan system limits, converting a theoretical measure into a usable engineering metric. Table 6 shows how to use this logic to distribute resources effectively in various situations.
Table 6: Expected total pipeline iterations for different LLM verification capabilities
LLM Capability Level Success Rate $\delta$ Expected Total Iterations $\mathbb{E}[n]$
@@
Furthermore, the exponential tail bound $\mathbb{P}(n>k)\leq\alpha(1-\delta)^{k}$ ensures that the probability of excessively long refinement loops decays exponentially. This enables engineers to configure reliable timeouts for safety-critical systems, ensuring that the verification pipeline fails gracefully rather than hanging indefinitely.